# Decoding Cortical Microcircuits: A Generative Model for Latent Space Exploration and Controlled Synthesis

### v1

@@ 5.3.1 SHAP Analysis
Using SHAP analysis shap to interpret feature contributions, we analyzed the relationship between the 32 latent dimensions and generated graph properties. Figure 3 illustrates the SHAP results for the clustering coefficient, showing that a few dimensions largely control this metric. Similar analyses were performed for each graph property (see Appendix 7.5).
Figure 3: SHAP analysis for clustering coefficient.
5.3.2 Discovering Key Latent Directions Affecting Generation Metrics
To visualize the main axes of topological structure variation across the microcircuits, we applied non-linear dimensionality reduction using t-SNE and identified topological features exhibiting gradients in the resulting embedding space (Figure 4). For each graph descriptor, we then aimed to find a consistent direction in the latent space that correlates with changes in its value. To achieve this, we divided the 500 test graphs into 20 value bins, each containing 5% of the data, and trained a linear regression model to predict the bin index from the 32-dimensional latent vector. The constant gradient direction of the linear regression model is advantageous for identifying this direction linear_reg , and further details of this method are in Appendix 7.6. We plot the $\text{R}^{2}$ scores of these linear regression models in Figure 5(a).
Since there might be some inherent correlation between the metrics, which results in similar or divergent gradient directions for different metrics, we computed the correlation between each pair of metrics directly from the test dataset, as shown in Figure 5(b). We also computed the cosine similarity between each pair of gradient directions, as shown in Figure 5(c). The relative magnitude of the cosine similarity between the correlation values and the gradient directions is largely consistent, further confirming that the gradient direction does indeed correspond to specific metrics. Further details regarding the correlations between different metrics are provided in Appendix7.8.
To validate whether these identified directions truly reflect changes in the corresponding metrics, we performed a traversal experiment. Starting from the mean latent vector of the test set, we moved along each metric’s identified gradient direction and decoded the resulting latent vectors to observe the change in the metric’s value. The details are shown in Appendix 7.7.
Refer to caption(a) Mean degreeRefer to caption(b) EfficiencyRefer to caption(c) TransitivityRefer to caption(d) AssortativityRefer to caption(e) Clus. Coef.Refer to caption(f) Modularity
@@ 5.3.2 Discovering Key Latent Directions Affecting Generation Metrics
Figure 4: t-SNE visualizations of vector embeddings reveal axes of variation in connectome topology. The darkness of each point’s color corresponds to the magnitude of its associated bin index for the specific metric being considered. To further aid visualization, the red line overlaid on the t-SNE plot represents the direction of metric variation, obtained by fitting an auxiliary 2D linear regression model to the 2D t-SNE embeddings. This 2D regression is solely for visual convenience and does not replace the 32-dimensional analysis described in the main text.
Refer to caption(a) $\text{R}^{2}$ scoresRefer to caption(b) CorrelationsRefer to caption(c) Gradient distance
Figure 5: (a) $\text{R}^{2}$ scores of the linear regression models predicting each of the six graph metrics from the 32-dimensional latent embeddings, indicating the predictive strength of the latent space. (b) Spearman’s rank correlation coefficient matrix between the six graph metric descriptors, calculated directly from the test dataset. (c) Cosine distance matrix between the gradient directions in the 32-dimensional latent space for each of the six graph metrics, illustrating the similarity in how different metrics are encoded in the latent space.
5.4 A Pipeline for Controllable Microcircuit Generation With MCMC Sampling
@@
For the investigation of structure-function relationships in brain networks, the ability to generate network samples that closely match specific target descriptor values would be invaluable. Here, we propose a general pipeline to achieve this goal, leveraging our latent VAE connectome generator in conjunction with Markov Chain Monte Carlo (MCMC) sampling MCMC .
@@
As direct integration to compute $p(\Omega_{\mathcal{T}})$ is not feasible, we utilize Markov Chain Monte Carlo (MCMC) MCMC to sample from the target conditional probability distribution despite the unknown denominator. We define a log-target-density function (LTD) to evaluate the likelihood of a proposed latent vector $\mathbf{z}$:
@@
Thus, the target conditional distribution for should be proportional to $\text{exp}(\text{LTD}(\mathbf{z}))\cdot\mathbb{I}(\mathbf{z}\in\Omega_{\mathcal{T}})$. We utilize the Metropolis-Hastings acceptance rule to generate a sequence of $N$ sampled latent vectors. The detailed steps of the sampling algorithm are provided in the Appendix 7.9.
To demonstrate the ability to generate graphs satisfying specific properties, we conducted experiments by specifying target percentile values ranging from 0% to 100%. We set $N=500$ in the experiment. The mean values of the graph metrics for the 500 generated graphs at different target percentiles are shown in Figure 6. Generated graph metric values that show a roughly monotonically increasing trend with the target bin index demonstrates the effectiveness of this method. Furthermore, examples of the generated graphs for different target metrics and target values are presented in the Appendix7.10.
Figure 7 presents examples of controlled graph generation where the mean degree is the targeted property, showing samples generated by setting different target percentile ranges for this metric. Detailed results for controlling other graph properties are provided in the Appendix 7.10.
@@
6 Conclusions and Discussions
In this work, we introduced a VAE-based generative model that successfully learns a compact and interpretable latent representation of mouse cortical microcircuit topology from detailed connectome data. We demonstrated that specific, interpretable directions within this latent space correspond to meaningful network properties, enabling the controlled synthesis of novel microcircuits with desired structural features via latent space navigation and an MCMC-based pipeline. This approach offers a powerful new tool to investigate the underlying design principles of neural circuits, explore structure-function relationships, and potentially inform the development of more advanced artificial neural networks by elucidating how complex structures might arise from low-dimensional generative rules.

### v2

@@ Introduction
Using SHAP analysis (Lundberg and Lee 2017) to interpret feature contributions, we analyzed the relationship between the 32 latent dimensions and generated graph properties. Figure 3 illustrates the SHAP results for the clustering coefficient, showing the entangled relationship between the latent dimensions and the graph metrics. Similar analyses were performed for each graph property (see Appendix E). These results motivated our subsequent development of a controllable generation method that operates effectively on such an entangled representation.
Figure 3: SHAP analysis for clustering coefficient. The results of SHAP analysis show the entangled relationship between the latent dimensions and the graph metrics.
Discovering Key Latent Directions Affecting Generation Metrics
We first visualized the microcircuit topology with t-SNE (Figure 4). To identify a latent direction for each graph metric, we trained a linear model to predict its value bin (discretized into 20 quantiles from 500 test graphs) from the 32D latent vectors. The resulting constant gradient defines each direction (Appendix F), and the model’s high $R^{2}$ scores confirm a strong fit (Figure 12, left).
We then verified that these directions correspond to distinct metrics. The pairwise correlations between the metrics themselves (Figure 12, middle) closely matched the cosine similarities of their corresponding latent directions (Figure 12, right), confirming their specificity (Appendix H).
Finally, a traversal experiment validated these directions. Moving along a specific gradient from a mean latent vector and then decoding controllably manipulated the corresponding metric in the generated circuits, confirming that the directions directly govern structural features (Appendix G).
(a) Mean degree
(b) Efficiency
(c) Transitivity
(d) Assortativity
(e) Clus. Coef.
(f) Modularity
@@
Figure 4: t-SNE visualizations of the latent space reveal that variations in specific topological metrics correspond to distinct gradient directions within the embeddings. The darkness of each point’s color corresponds to the magnitude of its associated bin index for the specific metric being considered. To further aid visualization, the red line overlaid on the t-SNE plot represents the direction of metric variation, obtained by fitting an auxiliary 2D linear regression model to the 2D t-SNE embeddings. This 2D regression is solely for visual convenience and does not replace the 32-dimensional analysis described in the main text.
A Pipeline for Controllable Microcircuit Generation With MCMC Sampling
@@
For the investigation of structure-function relationships in brain networks, the ability to generate network samples that closely match specific target descriptor values would be invaluable. Here, we propose a general pipeline to achieve this goal, leveraging our latent VAE connectome generator in conjunction with Markov Chain Monte Carlo (MCMC) sampling (Metropolis et al. 1953).
@@
As direct integration to compute $p(\Omega_{\mathcal{T}})$ is not feasible, we utilize Markov Chain Monte Carlo (MCMC) (Metropolis et al. 1953) to sample from the target conditional probability distribution despite the unknown denominator. We define a log-target-density function (LTD) to evaluate the likelihood of a proposed latent vector $\mathbf{z}$:
@@
Thus, the target conditional distribution for $\mathbf{z}$ should be proportional to $\text{exp}(\text{LTD}(\mathbf{z}))\cdot\mathbb{I}(\mathbf{z}\in\Omega_{\mathcal{T}})$. We utilize the Metropolis-Hastings acceptance rule to generate a sequence of $N$ sampled latent vectors. The detailed steps and parameters of the sampling algorithm are provided in the Appendix I.
To demonstrate the ability to generate graphs satisfying specific properties, we conducted experiments by specifying target percentile values ranging from 0% to 100%. We set $N=500$ in the experiment. Figure 5 shows how the mean values of four key graph metrics evolve with the target percentile; the complete set of six curves is deferred to the appendix (Figure 17) for brevity. Generated graph metric values that show a roughly monotonically increasing trend with the target bin index demonstrates the effectiveness of this method. Furthermore, examples of the generated graphs for different target metrics and target values are presented in the AppendixJ.
Figure 5: Metrics of generated graphs when setting different targets (4 of 6 target graph metrics).
Figure 6 presents examples of controlled graph generation where the mean degree is the targeted property, showing samples generated by setting different target percentile ranges for this metric. Detailed results for controlling other graph properties are provided in the Appendix J.
@@
Explore the Relationship between Structure and Function
Given that cortical microcircuits inherently operate as recurrent neural networks (RNNs), we employ reservoir computing (RC) to assess their functional capabilities. We use performance on classic tasks, specifically memory (copying) and classification, as a proxy for the network’s functional features. To investigate this, we constructed reservoir networks using connectome architectures from the VAE and tested them on copying and classification tasks (Appendix K). We primarily investigated two questions: 1. Do these connectome-based reservoirs outperform randomly connected ones? 2. Are there graph features that, when systematically varied, predictably modulate task-specific performance? All of the hyperparameters of the experiments are stated in Appendix K.
(a) Efficiency
(b) Mean Degree
(c) Transitivity
(d) Clustering Coefficient
Figure 7: 4 of the 6 metrics have a significant impact on task performance.
Copy Task
Following (Keller et al. 2024), we use a copy task to assess network memory retention (details in Appendix K.1). We specified target values for a set of features at five distinct percentiles (25%, 37.5%, 50%, 62.5%, 75%) of their distributions. For each condition, we constructed Reservoir Networks from three VAE-generated connectome graphs, yielding an average test loss $L_{gen}$. As a baseline, we repeated this process using density-matched random graphs to obtain a corresponding loss $L_{random}$. The performance gain of the VAE-generated structures over their random counterparts is quantified by the percentage decrease in test loss, $\delta$, calculated as follows:
$\delta=\displaystyle\frac{L_{random}-L_{gen}}{L_{random}}\times 100\%$ (4)
The results are documented in Table 2:
Target Percentile 25% 37.5% 50% 62.5% 75% Ave.
Mean Deg. 39.94 54.32 55.23 50.34 55.36 51.04
Clus. Coef. 52.67 52.69 52.45 53.45 49.03 52.06
Efficiency 47.64 59.29 59.31 54.61 52.00 54.57
Transitivity 52.23 45.53 56.05 58.93 51.15 52.78
Assortativity 49.68 57.55 52.01 45.28 53.71 51.64
Modularity 63.98 60.03 58.42 49.82 57.20 57.89
Table 2: Performance Advantage of VAE-Generated Connectomes over Random Baselines in Reservoir Tasks (Measured by % Test Loss Decrease)
Appendix K.1 visually compares the copy task outputs from networks with VAE-generated reservoirs against those with random counterparts.
Classification Task
On the sequential MNIST classification task, we found a clear monotonic or near-monotonic relationship between a reservoir network’s test accuracy and several of its reservoir’s graph features—notably efficiency, mean degree, transitivity, and clustering coefficient (Figure 7). This result was established by training networks built from reservoir graphs systematically generated to have feature values at five distinct percentiles (25-75%). Specifically, for each percentile condition, we randomly selected three generated graphs, constructed an network from each, and averaged their final test accuracies (details in Appendix K.2).
Conclusions and Discussions
We introduced a VAE-based generative model that learns a compact, interpretable latent space for mouse cortical microcircuit topology. We demonstrated that directions in this space encode specific network properties, enabling the controlled synthesis of novel circuits with desired features via latent navigation and an MCMC pipeline. This approach provides a powerful tool to investigate neural design principles, explore structure-function relationships, and inform advanced AI development by revealing the low-dimensional generative rules behind complex structures.